PSa 2-10 Calculate Gross Pay for a Tipped Employee

Calculate gross pay for each of the following employees. All are paid an overtime wage rate that is 1.5 times their respective regular wage rates.

NOTE: For simplicity, all calculations throughout this exercise, both intermediate and final, should be rounded to two decimal places at each calculation.

1: Anita Workman receives tips from customers as a standard component of her weekly pay. She is paid $2.50/hour by her employer and receives $284 in tips during the most recent 46-hour workweek.

Gross Pay = $

2: Cole Earnhardt receives tips from customers as a standard component of his weekly pay. He is paid 2.13/hour by his employer and receives $442 in tips during the most recent 42-hour workweek.

Gross Pay = $

3: Calista Flood receives tips from customers as a standard component of her weekly pay. She is paid $4.10/hour by her employer and receives $350 in tips during the most recent 38-hour workweek.

Gross Pay = $

4: Bethany Pugh receives tips from customers as a standard component of her weekly pay. She is paid $3.60/hour by her employer and receives $162 in tips during the most recent 51-hour workweek.

Gross Pay = $
Answer #1
Sr. No. . Name of Employee Particulars

Minimum Pay

(No. of hrs. X rate/hr)

Overtime Pay

(no. of OT hrs X rate /hr)

Gross Pay

(Minimum Pay + OT Pay

Tipped

Wages

No. of tips X rate

1 Anita OT pay=2.5 X 1.5=3.75 40 hrs X $.2.50/hr = $.100 6 hrs. X $ 3.75/hr = $. 22.5 $ 122.5 284 X (5.12-2.50) $.2.62 = $. 744.08
2. Cole Earnhardt OT pay=2.13 X 1.5=3.20 40 hrs X $.2.13/hr = $.85.2 2 hrs. X $ 3.20/hr = $. 6.40 $ 91.6 442 X (5.12-2.99) $.2.99 = $. 1321.58
3 Calista Flood 38 hrs X $.4.10/hr = $.155.8 -------- $ 155.8 350 X (5.12-4.10) $.1.02 = $. 357
4 Bethany Pugh OT pay=3.6 X 1.5=5.4 40 hrs X $.3.6/hr = $.144 11 hrs. X $ 5.4/hr = $. 59.40 $ 203.4 162 X (5.12-3.6) $.1.52 = $. 246.24
Answer #2

Anita Workman receives tips from customers as a standard component of her weekly pay. She is paid $2.50/hour by her employer and receives $284 in tips during the most recent 46-hour workweek

Step A: 2.50(Below minimum wage) x 40 = 100(pay at regular rate)

Step B: 7.25(actual min. wage) - 2.50= 4.75 Tip Credit

Step C: 7.25 x 1.5= 10.88

Step D: 10.88 - 4.75 = 6.13

Step E:  6.13 x 6 = 36.78 overtime pay 

Step F: 100regular pay + 36.78overtime pay+ 284Tips=420.78 gross pay

source: eLab
answered by: J.Louis